# CNN-JEPA: Self-Supervised Pretraining Convolutional Neural Networks Using Joint Embedding Predictive Architecture

### v1

@@ V-A Comparison to other methods
V-B Masking strategy
TABLE II: Ablation on masking strategies on ImageNet-100.
Linear k-NN
Masking strategy top-1 top-1
Mixed 73.26 56.26
Multi-block 73.02 55.62
Random 57.64 36.64
We experiment with different masking strategies and present our results in table II. Random masking, common in most masked autoencoding methods (e.g. [2, 1, 3]), performs poorly, which is in line with the findings of I-JEPA [4]. The multi-block masking strategy proposed by Bojanowski et al. [4] performs well on convolutional networks as well. Additionally, we propose a mixed masking strategy, where we randomly select between the multi-block and random masking strategies for each minibatch. To primarily use multi-block masking, we choose the probability of using multi-block masking to be 0.75. As shown in table II, the mixed masking strategy performs better than both random and multi-block masking, outperforming the latter by a small margin both on linear and k-NN classification.
V-C Predictor architecture
TABLE III: Ablation on using depthwise separable convolution in the predictor on ImageNet-100.
Predictor Linear k-NN Predictor Training cost
convolution top-1 top-1 size [M] (GPU hours)
Standard 72.64 54.76 12.7 M 62.07
Depthwise separable 73.26 56.26 113 M 51.51
TABLE IV: Predictor depth ablation on ImageNet-100 and ImageNet-1k.
Predictor Linear k-NN Predictor
Dataset # layers top-1 top-1 size [M]
ImageNet-100 1 64.86 51.84 4.2 M
2 73.22 59.22 8.4 M
3 73.26 56.26 12.7 M
4 71.72 53.84 16.9 M
5 71.14 50.94 21.1 M
6 69.24 47.68 25.3 M
ImageNet-1k 2 49.78 31.96 8.4 M
3 54.23 35.00 12.7 M
TABLE V: Predictor kernel size ablation on ImageNet-100.
Predictor Linear k-NN Predictor
kernel size top-1 top-1 size [M]
3 73.26 56.26 12.7 M
5 73.08 54.88 12.8 M
7 72.44 53.76 12.9 M
9 73.16 53.46 13.1 M
We conduct ablation on the predictor architecture and present the results in tables V, IV and III. Table III clearly shows the benefits of using depthwise separable convolution in the predictor. Depthwise separable convolution not only requires 90% fewer parameters than standard convolution and reduces training time by 17%, but also improves the linear and k-NN top-1 accuracies by 0.62% and 1.5% respectively on ImageNet-100.
The depth of the predictor also has significant impact on pretraining performance. As shown in table IV, two or three layers are optimal, even though such predictors have a smaller receptive field than deeper predictors. On ImageNet-100, linear and k-NN top-1 accuracies indicate different optimal depths; however, on ImageNet-1k, the three-layer predictor provides the highest accuracy by a large margin, 6% and 3% respectively.
Table V shows our results on the predictor’s kernel size. We find that 3x3 kernels provide the highest accuracies; however, the difference between all tested kernel sizes is less than 3%.
@@ VI Conclusions
In this paper, we present CNN-JEPA, a novel self-supervised learning method that successfully adapts I-JEPA to convolutional networks. Our method overcomes the challenges of applying masked inputs to CNNs through the use of a sparse CNN encoder and carefully designed masking strategy. To efficiently handle the joint embedding prediction task, our method utilizes a fully convolutional predictor with depthwise separable convolutions. We demonstrated that CNN-JEPA outperforms I-JEPA with ViT architectures on ImageNet-100 and is competitive with other CNN-based SSL methods on both ImageNet-100 and ImageNet-1k, achieving top-1 linear accuracies of 73.26% and 54.23% respectively. Our proposed mixed masking strategy, combining multi-block and random masking, showed improved performance over either strategy alone. Ablation studies on the predictor architecture revealed the benefits of using depthwise separable convolutions and optimal depth setting for the predictor. CNN-JEPA provides a simpler and more efficient approach to self-supervised learning for CNNs, requiring minimal augmentations and no separate projector network.
@@ Acknowledgements
ChatGPT‑4o and Claude Sonet 3.5 were used to improve the quality of the text in this paper.
We acknowledge KIFÜ (Governmental Agency for IT Development, Hungary, https://ror.org/01s0v4q65) for awarding us access to the Komondor HPC facility based in Hungary.

### v2

@@ VI Conclusions
In this paper, we present CNN-JEPA, a novel self-supervised learning method that successfully adapts I-JEPA to convolutional networks. Our method overcomes the challenges of applying masked inputs to CNNs through the use of a sparse CNN encoder and carefully designed masking strategy. To efficiently handle the joint embedding prediction task, our method utilizes a fully convolutional predictor with depthwise separable convolutions. We demonstrated that CNN-JEPA outperforms I-JEPA with ViT architectures on ImageNet-100 and is competitive with other CNN-based SSL methods on both ImageNet-100 and ImageNet-1k, achieving top-1 linear accuracies of 73.26% and 54.23% respectively. CNN-JEPA provides a simpler and more efficient approach to self-supervised learning for CNNs, requiring minimal augmentations and no separate projector network.
@@ Acknowledgements
The work reported in this paper carried out at BME, has been supported by the European Union project RRF-2.3.1-21-2022-00004 within the framework of the Artificial Intelligence National Laboratory. Project no. TKP2021-NVA-02 and 149336 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA and MEC_R_24 funding schemes respectively. This research was partly funded by the National Research, Development and Innovation Office of Hungary (FK 142163 grant). We gratefully acknowledge the support of NVIDIA with the donation of the A100 GPU used for this research and KIFÜ (Governmental Agency for IT Development, Hungary, https://ror.org/01s0v4q65) for awarding us access to the Komondor HPC facility based in Hungary.